OpenStudy (anonymous):

solve with a matrix equation: 4x-3y=15 2x+y=5

OpenStudy (anonymous):

\[ \left[\begin{matrix}4&-3\\2&1\end{matrix}\right]\cdot \left[\begin{matrix}x\\y\end{matrix}\right]= \left[\begin{matrix}15\\5\end{matrix}\right]\\ A\cdot X=B\\ X=A^{-1}B \] find A inverse and multiply with B